# Hierarchical certification of non-classical network correlations

### v3

@@ 3 Nonlinear $\ell$-(Q)NN hierarchies for star networks
4 Linear FNN inequalities for chains
Up to now, we have used nonlinear inequalities to detect $\ell$-NN and $\ell$-QNN. On one hand, nonlinear inequalities are useful because they approximate better the boundaries of the relevant sets, which are known to be non-linear [25]. However, the nonlinearities make it difficult to optimize these inequalities and to find explicit models. In this section and in the following one we present linear inequalities whose violations detect FNN. For deriving them, we exploit the fact that realizing entanglement swapping is not possible when parties are connected by classical sources. It is known that at least some supra-quantum theories do not allow for entanglement swapping [43], and this fact has been exploited to develop NN inequalities in the tripartite chain scenario [44]. However, we are interested in guaranteeing non-classical phenomena, so as long as there exists a non-classical theory that allows for entanglement swapping (as it is the case of quantum mechanics), the inequalities below can, in principle, be used for detecting FNN.
4.1 The tripartite chain
Let us begin, as in Section 2, with analyzing the tripartite chain scenario, depicted in Fig. 1 (a). A way to guarantee that both sources are not classical is by completing entanglement swapping, so that after Bob’s measurement the joint state of Alice and Charlie can violate the CHSH inequality. Using this fact, consider the following operator, where Alice can perform two dichotomic measurements, Bob performs a single, four-outcome measurement, and Charlie can perform three dichotomic measurements:
@@ 4.1 The tripartite chain
$\mathcal{B}_{3}=\,$ $C_{0}(B^{0}A_{0}+B^{0}A_{1}+B^{1}A_{0}+B^{1}A_{1})$
$+C_{1}(B^{0}A_{0}-B^{0}A_{1}-B^{1}A_{0}+B^{1}A_{1})$
$+C_{1}(B^{2}A_{0}-B^{2}A_{1}-B^{3}A_{0}+B^{3}A_{1})$
$+C_{2}(B^{2}A_{0}+B^{2}A_{1}+B^{3}A_{0}+B^{3}A_{1})$
@@
±$\qquad\,\,\pm[2(B^{0}-B^{3})A_{1}+2(B^{1}-B^{2})A_{0}],\pm[2(B^{3}-B^{0})A_{0}+2(B^{2}-B^{1})A_{1}]\}$
@@
where the last bound can be found by individually obtaining the maximum value for each case, using $|A_{i}|\leq 1$ and $\sum_{i=0}^{3}B^{i}=\mathbbm{1}$.
@@
$\mathcal{B}_{3}=\,$ $A_{0}(B^{0}C_{0}+B^{1}C_{0}+B^{0}C_{1}-B^{1}C_{1})$
$+A_{1}(B^{0}C_{0}+B^{1}C_{0}-B^{0}C_{1}+B^{1}C_{1})$
$+A_{0}(B^{2}C_{1}-B^{3}C_{1}+B^{2}C_{2}+B^{3}C_{2})$
$+A_{1}(-B^{2}C_{1}+B^{3}C_{1}+B^{2}C_{2}+B^{3}C_{2})$
@@
$A_{0}B^{0}C_{0}+A_{0}B^{1}C_{0}+A_{0}B^{0}C_{1}-A_{0}B^{1}C_{1}$ (24)
$+A_{1}B^{0}C_{0}+A_{1}B^{1}C_{0}-A_{1}B^{0}C_{1}+A_{1}B^{1}C_{1}$
$+A_{0}B^{2}C_{1}-A_{0}B^{3}C_{1}+A_{0}B^{2}C_{2}+A_{0}B^{3}C_{2}$
$-A_{1}B^{2}C_{1}+A_{1}B^{3}C_{1}+A_{1}B^{2}C_{2}+A_{1}B^{3}C_{2}\leq 2$
@@
4.2 Long chains
The reasoning above can be generalized to chain networks of arbitrary length, by exploiting the fact that non-classical correlations between the extremal parties can only be established if all the sources are non-classical. Consider thus the following operator:
$\mathcal{B}_{n}:=$ $A_{0}^{(1)}B^{00}A_{0}^{(n)}+A_{0}^{(1)}B^{01}A_{0}^{(n)}+A_{0}^{(1)}B^{00}A_{1}^{(n)}-A_{0}^{(1)}B^{01}A_{1}^{(n)}$ (25)
$+A_{1}^{(1)}B^{00}A_{0}^{(n)}+A_{1}^{(1)}B^{01}A_{0}^{(n)}-A_{1}^{(1)}B^{00}A_{1}^{(n)}+A_{1}^{(1)}B^{01}A_{1}^{(n)}$
$+A_{0}^{(1)}B^{10}A_{1}^{(n)}-A_{0}^{(1)}B^{11}A_{1}^{(n)}+A_{0}^{(1)}B^{10}A_{2}^{(n)}+A_{0}^{(1)}B^{11}A_{2}^{(n)}$
$-A_{1}^{(1)}B^{10}A_{1}^{(n)}+A_{1}^{(1)}B^{11}A_{1}^{(n)}+A_{1}^{(1)}B^{10}A_{2}^{(n)}+A_{1}^{(1)}B^{11}A_{2}^{(n)},$
where the extremal parties are $\textsf{A}_{1}$ and $\textsf{A}_{n}$, which have operators $A^{(1)}_{i}$ and $A^{(n)}_{i}$, respectively, and $B^{st}$ are products of operators of all the non-extremal parties, given by $B^{st}=\sum_{{}^{\oplus_{k=2}^{n-1}i_{k}j_{k}=st}}M^{i_{2}j_{2}}\cdots M^{i_{n-1}j_{n-1}}$, with $\oplus$ denoting the bitwise XOR operation, i.e., $xy\oplus uv=(x\oplus u,y\oplus v)$, and $M^{i_{k}j_{k}}$ are the projectors corresponding to the single measurement of party $\textsf{A}_{k}$. Similar to Eq. (24), the extremal party $\textsf{A}_{n}$ has three dichotomic measurements.
@@ 4.2 Long chains
$\mathcal{B}_{n}:=\,$ $A_{0}^{(1)}(B^{00}A_{0}^{(n)}+B^{01}A_{0}^{(n)}+B^{00}A_{1}^{(n)}-B^{01}A_{1}^{(n)})$
$+A_{1}^{(1)}(B^{00}A_{0}^{(n)}+B^{01}A_{0}^{(n)}-B^{00}A_{1}^{(n)}+B^{01}A_{1}^{(n)})$
$+A_{0}^{(1)}(B^{10}A_{1}^{(n)}-B^{11}A_{1}^{(n)}+B^{10}A_{2}^{(n)}+B^{11}A_{2}^{(n)})$
$+A_{1}^{(1)}(-B^{10}A_{1}^{(n)}+B^{11}A_{1}^{(n)}+B^{10}A_{2}^{(n)}+B^{11}A_{2}^{(n)})$
[…]
±$\qquad\,\,\pm[2(B^{00}-B^{01})A_{1}^{(n)}+2(B^{10}-B^{11})A_{2}^{(n)}]\}$
[…]
where the terms in the maximization are found by applying the triangle inequality and forming groups that can be jointly maximized. Then, the bound follows from optimizing individually each of the elements under the constraints $\sum_{b_{0},b_{1}}B^{b_{0}b_{1}}=\mathbbm{1}$ and $|A_{i}^{(n)}|\leq 1$.
Now, consider the case where the classical source is not in the extreme of the chain. In that case, we can re-write $\mathcal{B}_{n}$ as
$\mathcal{B}_{n}=\,$ $B^{00}\left(A_{0}^{(1)}A_{0}^{(n)}+A_{0}^{(1)}A_{1}^{(n)}+A_{1}^{(1)}A_{0}^{(n)}-A_{1}^{(1)}A_{1}^{(n)}\right)$
$+B^{01}\left(A_{0}^{(1)}A_{0}^{(n)}-A_{0}^{(1)}A_{1}^{(n)}+A_{1}^{(1)}A_{0}^{(n)}+A_{1}^{(1)}A_{1}^{(n)}\right)$
$+B^{10}\left(A_{0}^{(1)}A_{1}^{(n)}-A_{0}^{(1)}A_{2}^{(n)}+A_{1}^{(1)}A_{1}^{(n)}+A_{1}^{(1)}A_{2}^{(n)}\right)$
$+B^{11}\left(-A_{0}^{(1)}A_{1}^{(n)}+A_{0}^{(1)}A_{2}^{(n)}+A_{1}^{(1)}A_{1}^{(n)}+A_{1}^{(1)}A_{2}^{(n)}\right).$
Each line above corresponds to a different CHSH test, weighted by a different probability of obtaining certain outcomes in the central parties (recall that the $B^{st}$ are defined by $B^{st}=\sum_{\oplus_{\ell=2}^{n-1}i_{\ell}{}j_{\ell}=st}M^{i_{2}j_{2}}\dots M^{i_{n-1}j_{n-1}}$, with $\oplus$ being the bit-wise XOR). Therefore, each of the lines above can be bounded by $2\sum_{\oplus_{i=2}^{n-1}a_{i}=00}p(a_{2},\dots,a_{n-1})$, and using the fact that $\sum_{i=2}^{n-1}\sum_{a_{i}}p(a_{2},\dots,a_{n-1})=1$ the upper bound of $2$ follows. We are thus left with the inequality
$\mathcal{B}_{n}=\,$ $A_{0}^{(1)}\left[(B^{00}+B^{01})A_{0}^{(n)}+(B^{00}-B^{01})A_{1}^{(n)}\right]$ (26)
$+A_{1}^{(1)}\left[(B^{00}+B^{01})A_{0}^{(n)}+(B^{00}-B^{01})A_{1}^{(n)}\right]$
$+A_{0}^{(1)}\left[(B^{10}-B^{11})A_{1}^{(n)}+(B^{10}+B^{11})A_{2}^{(n)}\right]$
$+A_{1}^{(1)}\left[(-B^{10}+B^{11})A_{1}^{(n)}+(B^{10}+B^{11})A_{2}^{(n)}\right]$
[…]
@@
5 Linear FNN inequalities for stars
It is possible to use the results of the previous section to witness FNN in star networks, in two different ways. The first one is by decomposing the star network into a collection of tripartite chain subnetworks comprised by the parties $(\textsf{A}_{i},\textsf{A}_{j\not=i},\textsf{B})$, and testing Eq. (24) on each of them. This leads to witnesses in scenarios where the central party performs a single, 4-outcome measurement, and the branch parties can perform one of either two or three measurements, depending on the particular inequalities that are used for the testing.
The second one is by exploiting the same principle, namely that only measurements on non-classical systems can produce non-local correlations. In this case, we are interested in detecting genuinely multipartite nonlocality in the branch parties after the measurement of the central party. One way of detecting genuine multipartite nonlocality for $n$ parties is via the Mermin-Svetlichny inequality [9, 45, 46]:
$\mathcal{S}(A_{x_{1}}^{(\vec{i})},\dots,A_{x_{n}}^{(\vec{i})})\leq 2^{n-1},$ (27)
where the operator $\mathcal{S}(A_{1},\dots,A_{n})\equiv S_{n}$ is built recursively via
@@ 5 Linear FNN inequalities for stars
$\sum_{\vec{i}}M^{\vec{i}}\mathcal{S}^{(\vec{i})}(A_{x_{1}}^{(\vec{i})},\dots,A_{x_{n}}^{(\vec{i})})\leq 2^{n-1},$ (28)
@@
$\gamma_{0i_{[n-1]}}=\frac{1}{\sqrt{p(\vec{i})}}\alpha_{0}\prod_{j=2}^{n}\alpha_{i_{j}},\qquad\delta_{1i_{[n-1]}}=\frac{1}{\sqrt{p(\vec{i})}}\alpha_{1}\prod_{j=2}^{n}\alpha_{\overline{i}_{j}},$
@@
6 Inequalities for general networks

### v4

@@ IV Nonlinear $\ell$-(Q)NN hierarchies for star networks
V Linear FNN inequalities for chains
Up to now, we have used nonlinear inequalities to detect $\ell$-NN and $\ell$-QNN. On one hand, nonlinear inequalities are useful because they approximate better the boundaries of the relevant sets, which are known to be nonlinear [25]. However, the nonlinearities make it difficult to optimize these inequalities and to find explicit models. In this section and in the following one we present linear inequalities whose violations detect FNN. For deriving them, we exploit the fact that realizing entanglement swapping is not possible when parties are connected by classical sources. It is known that at least some supra-quantum theories do not allow for entanglement swapping [43], and this fact has been exploited to develop NN inequalities in the tripartite chain scenario [44]. However, we are interested in guaranteeing nonclassical phenomena, so as long as there exists a nonclassical theory that allows for entanglement swapping (as it is the case of quantum mechanics), the inequalities below can, in principle, be used for detecting FNN.
V.1 The tripartite chain
Let us begin, as in Section III, with analyzing the tripartite chain scenario, depicted in Fig. 1 (a). A way to guarantee that both sources are not classical is by completing entanglement swapping, so that after Bob’s measurement the joint state of Alice and Charlie can violate the CHSH inequality. Using this fact, consider the following operator, where Alice can perform two dichotomic measurements, Bob performs a single, four-outcome measurement, and Charlie can perform three dichotomic measurements:
@@ V.1 The tripartite chain
$\left\langle\mathcal{B}_{3}\right\rangle=\,$ $\left\langle C_{0}(B^{0}A_{0}+B^{0}A_{1}+B^{1}A_{0}+B^{1}A_{1})\right\rangle$
$+\left\langle C_{1}(B^{0}A_{0}-B^{0}A_{1}-B^{1}A_{0}+B^{1}A_{1})\right\rangle$
$+\left\langle C_{1}(B^{2}A_{0}-B^{2}A_{1}-B^{3}A_{0}+B^{3}A_{1})\right\rangle$
$+\left\langle C_{2}(B^{2}A_{0}+B^{2}A_{1}+B^{3}A_{0}+B^{3}A_{1})\right\rangle$
@@
±$\pm\langle 2(B^{0}+B^{2})A_{1}+2(B^{1}+B^{3})A_{0}\rangle,$
±$\pm\langle 2(B^{0}-B^{3})A_{1}+2(B^{1}-B^{2})A_{0}\rangle,$
±$\pm\langle 2(B^{3}-B^{0})A_{0}+2(B^{2}-B^{1})A_{1}\rangle\}$
@@
where the last bound can be found by individually obtaining the maximum value for each case, using $|\left\langle A_{i}\right\rangle|\leq 1$ and $𝟙\sum_{i=0}^{3}B^{i}=\openone$.
@@
$\left\langle\mathcal{B}_{3}\right\rangle=\,$ $\left\langle A_{0}(B^{0}C_{0}+B^{1}C_{0}+B^{0}C_{1}-B^{1}C_{1})\right\rangle$
$+\left\langle A_{1}(B^{0}C_{0}+B^{1}C_{0}-B^{0}C_{1}+B^{1}C_{1})\right\rangle$
$+\left\langle A_{0}(B^{2}C_{1}-B^{3}C_{1}+B^{2}C_{2}+B^{3}C_{2})\right\rangle$
$+\left\langle A_{1}(-B^{2}C_{1}+B^{3}C_{1}+B^{2}C_{2}+B^{3}C_{2})\right\rangle$
@@
±$\qquad\,\,\,\left.\pm\left\langle 2(B^{0}-B^{1})C_{1}+2(B^{2}-B^{3})C_{2}\right\rangle\right\}$
@@
$\left\langle A_{0}B^{0}C_{0}+A_{0}B^{1}C_{0}+A_{0}B^{0}C_{1}-A_{0}B^{1}C_{1}\right\rangle$ (21)
$+\left\langle A_{1}B^{0}C_{0}+A_{1}B^{1}C_{0}-A_{1}B^{0}C_{1}+A_{1}B^{1}C_{1}\right\rangle$
$+\left\langle A_{0}B^{2}C_{1}-A_{0}B^{3}C_{1}+A_{0}B^{2}C_{2}+A_{0}B^{3}C_{2}\right\rangle$
$-\left\langle A_{1}B^{2}C_{1}+A_{1}B^{3}C_{1}+A_{1}B^{2}C_{2}+A_{1}B^{3}C_{2}\right\rangle\leq 2$
@@
V.2 Long chains
The reasoning above can be generalized to chain networks of arbitrary length, by exploiting the fact that nonclassical correlations between the extremal parties can only be established if all the sources are nonclassical. Consider thus the following operator:
$\mathcal{B}_{n}:=$ $A_{0}^{(1)}B^{00}A_{0}^{(n)}+A_{0}^{(1)}B^{01}A_{0}^{(n)}+A_{0}^{(1)}B^{00}A_{1}^{(n)}$
$-A_{0}^{(1)}B^{01}A_{1}^{(n)}+A_{1}^{(1)}B^{00}A_{0}^{(n)}+A_{1}^{(1)}B^{01}A_{0}^{(n)}$
$-A_{1}^{(1)}B^{00}A_{1}^{(n)}+A_{1}^{(1)}B^{01}A_{1}^{(n)}+A_{0}^{(1)}B^{10}A_{1}^{(n)}$
$-A_{0}^{(1)}B^{11}A_{1}^{(n)}+A_{0}^{(1)}B^{10}A_{2}^{(n)}+A_{0}^{(1)}B^{11}A_{2}^{(n)}$
$-A_{1}^{(1)}B^{10}A_{1}^{(n)}+A_{1}^{(1)}B^{11}A_{1}^{(n)}+A_{1}^{(1)}B^{10}A_{2}^{(n)}$
$+A_{1}^{(1)}B^{11}A_{2}^{(n)},$
where the extremal parties are $\textsf{A}_{1}$ and $\textsf{A}_{n}$, which have operators $A^{(1)}_{i}$ and $A^{(n)}_{i}$, respectively, and $B^{st}$ are products of operators of all the non-extremal parties $\textsf{A}_{2},\cdots,\textsf{A}_{n-1}$, given by $B^{st}=\sum_{{}^{\oplus_{k=2}^{n-1}i_{k}j_{k}=st}}M^{i_{2}j_{2}}\cdots M^{i_{n-1}j_{n-1}}$, with $\oplus$ denoting the bitwise XOR operation, i.e., $xy\oplus uv=(x\oplus u,y\oplus v)$, and $M^{i_{k}j_{k}}$ are the projectors corresponding to the single measurement of party $\textsf{A}_{k}$. Similarly to Eq. (21), the extremal party $\textsf{A}_{n}$ has three dichotomic measurements.
@@ V.2 Long chains
$\left\langle\mathcal{B}_{n}\right\rangle$
$\quad=\left\langle A_{0}^{(1)}(B^{00}A_{0}^{(n)}\!+\!B^{01}A_{0}^{(n)}\!+\!B^{00}A_{1}^{(n)}\!-\!B^{01}A_{1}^{(n)})\right\rangle$
$\qquad+\left\langle A_{1}^{(1)}(B^{00}A_{0}^{(n)}\!+\!B^{01}A_{0}^{(n)}\!-\!B^{00}A_{1}^{(n)}\!+\!B^{01}A_{1}^{(n)})\right\rangle$
$\qquad+\left\langle A_{0}^{(1)}(B^{10}A_{1}^{(n)}\!-\!B^{11}A_{1}^{(n)}\!+\!B^{10}A_{2}^{(n)}\!+\!B^{11}A_{2}^{(n)})\right\rangle$
$\qquad+\left\langle A_{1}^{(1)}(\!-\!B^{10}A_{1}^{(n)}\!+\!B^{11}A_{1}^{(n)}\!+\!B^{10}A_{2}^{(n)}\!+\!B^{11}A_{2}^{(n)})\right\rangle$
$\quad\leq\,\max\left\{\pm\!\left\langle 2(B^{00}+B^{01})A_{0}^{(n)}+2(B^{10}+B^{11})A_{2}^{(n)}\right\rangle\!,\right.$
±$\qquad\qquad\quad\pm\!\!\left.\left\langle 2(B^{00}-B^{01})A_{1}^{(n)}+2(B^{10}-B^{11})A_{2}^{(n)}\right\rangle\right\}$
 $\quad\leq\,2,$
where the terms in the maximization are found by applying the triangle inequality and forming groups that can be jointly maximized. Then, the bound follows from optimizing individually each of the elements under the constraints $𝟙\sum_{b_{0},b_{1}}B^{b_{0}b_{1}}=\openone$ and $\left|\left\langle A_{i}^{(n)}\right\rangle\right|\leq 1$.
Now, consider the case where the classical source is not in the extreme of the chain. In that case, we can rewrite $\mathcal{B}_{n}$ as
$\mathcal{B}_{n}=$
$\quad B^{00}\left(A_{0}^{(1)}A_{0}^{(n)}+A_{0}^{(1)}A_{1}^{(n)}+A_{1}^{(1)}A_{0}^{(n)}-A_{1}^{(1)}A_{1}^{(n)}\right)$
$\qquad+B^{01}\left(A_{0}^{(1)}A_{0}^{(n)}\!-\!A_{0}^{(1)}A_{1}^{(n)}\!+\!A_{1}^{(1)}A_{0}^{(n)}\!+\!A_{1}^{(1)}A_{1}^{(n)}\right)$
$\qquad+B^{10}\left(A_{0}^{(1)}A_{1}^{(n)}\!-\!A_{0}^{(1)}A_{2}^{(n)}\!+\!A_{1}^{(1)}A_{1}^{(n)}\!+\!A_{1}^{(1)}A_{2}^{(n)}\right)$
$\qquad+B^{11}\left(\!-\!A_{0}^{(1)}A_{1}^{(n)}\!+\!A_{0}^{(1)}A_{2}^{(n)}\!+\!A_{1}^{(1)}A_{1}^{(n)}\!+\!A_{1}^{(1)}A_{2}^{(n)}\right)\!.$
Each line above corresponds to a different CHSH test, weighted by a different probability of obtaining certain outcomes in the central parties (recall that the $B^{st}$ are defined by $B^{st}=\sum_{\oplus_{\ell=2}^{n-1}i_{\ell}{}j_{\ell}=st}M^{i_{2}j_{2}}\dots M^{i_{n-1}j_{n-1}}$, with $\oplus$ being the bit-wise XOR). Therefore, the expectation value of each of the lines above can be bounded by $2\sum_{\oplus_{i=2}^{n-1}a_{i}=00}p(a_{2},\dots,a_{n-1})$, and using the fact that $\sum_{i=2}^{n-1}\sum_{a_{i}}p(a_{2},\dots,a_{n-1})=1$ the upper bound of $2$ follows. We are thus left with the inequality
$\left\langle\mathcal{B}_{n}\right\rangle$ $=\left\langle A_{0}^{(1)}\left[(B^{00}+B^{01})A_{0}^{(n)}+(B^{00}-B^{01})A_{1}^{(n)}\right]\right\rangle$ (22)
$\quad+\left\langle A_{1}^{(1)}\left[(B^{00}+B^{01})A_{0}^{(n)}+(B^{00}-B^{01})A_{1}^{(n)}\right]\right\rangle$
$\quad+\left\langle A_{0}^{(1)}\left[(B^{10}-B^{11})A_{1}^{(n)}+(B^{10}+B^{11})A_{2}^{(n)}\right]\right\rangle$
$\quad+\left\langle A_{1}^{(1)}\left[(-B^{10}+B^{11})A_{1}^{(n)}\!+\!(B^{10}+B^{11})A_{2}^{(n)}\right]\right\rangle$
$\leq 2.$
@@
VI Linear FNN inequalities for stars
It is possible to use the results of the previous section to witness FNN in star networks, in two different ways. The first one is by decomposing the star network into a collection of tripartite chain subnetworks comprised by the parties $(\textsf{A}_{i},\textsf{A}_{j\not=i},\textsf{B})$, and testing Eq. (21) on each of them. This leads to witnesses in scenarios where the central party performs a single, four-outcome measurement, and the branch parties can perform one of either two or three measurements, depending on the particular inequalities that are used for the testing.
The second one is by exploiting the same principle, namely that only measurements on nonclassical systems can produce nonlocal correlations. In this case, we are interested in detecting genuinely multipartite nonlocality in the branch parties after the measurement of the central party. One way of detecting genuine multipartite nonlocality for $n$ parties is via the Mermin-Svetlichny inequality [9, 45, 46]:
$\left\langle\mathcal{S}\left(A_{x_{1}}^{(\vec{i})},\dots,A_{x_{n}}^{(\vec{i})}\right)\right\rangle\leq 2^{n-1},$ (23)
where the operator $\mathcal{S}\left(A_{1},\dots,A_{n}\right)\equiv S_{n}$ is built recursively via
@@ VI Linear FNN inequalities for stars
$\left\langle\sum_{\vec{i}}M^{\vec{i}}\mathcal{S}^{(\vec{i})}\left(A_{x_{1}}^{(\vec{i})},\dots,A_{x_{n}}^{(\vec{i})}\right)\right\rangle\leq 2^{n-1},$ (24)
@@
$\gamma_{0i_{[n-1]}}=\frac{1}{\sqrt{p(\vec{i})}}\alpha_{0}\prod_{j=2}^{n}\alpha_{i_{j}},\quad\delta_{1i_{[n-1]}}=\frac{1}{\sqrt{p(\vec{i})}}\alpha_{1}\prod_{j=2}^{n}\alpha_{\overline{i}_{j}},$
@@
VII Inequalities for general networks